Suppose a star with radius 8.44 x 10^8 m has a peak wavelength of 677 nm in the spectrum of its emitted radiation.

(a) Find the energy of a photon with this wavelength.

J/photon

(b) What is the surface temperature of the star?

K

(c) At what rate is energy emitted from the star in the form of radiation? Assume the star is a blackbody (e = 1).

W

(d) Using the answer to part (a), estimate the rate at which photons leave the surface of the star.

photons/s
Answer #1

(A) energy of photon = hc/

= 6.626*10^-34*3*10^8 / 677*10^-9

= 2.93*10^-19 J/photon

(b) using wein's displacement law we have

max *T = 2898*10^-6 m

677*10^-9*T = 2898*10^-6

T= 4280.65 K

(C) rate at which energy is radiated =  AT^4

= 4pi*r^2T^4

= 5.67*10^-8*4pi*(8.44*10^8)^2 * ( 4280.65)^4

= 1.7*10^26 W

(D) no. of photons per second = 1.7*10^26 / 2.93*10^-19

= 5.8*10^44